OpenStudy (lovely_lovely):

Which expression can be used to find the area of a circle with a diameter of 40 cm? A. π · 20 B. π · 40 C. π · 202 D. π · 402

OpenStudy (knov):

area=pi*radius^2 is the formula.

OpenStudy (anonymous):

Okay, so before we answer this question we should probably understand the formula to find the area of a circle. I'll cheat for ya', it's r^2 x pi. The diameter is twice the radius, so the radius is 20. 20^2 = 400. I'd think the answer is 400 x pi but... o.O

OpenStudy (anonymous):

Are you sure D. isn't 400?

OpenStudy (anonymous):

I don't think it's any of them...

OpenStudy (knov):

i thinks the answers are not correctly written.

OpenStudy (anonymous):

That's what I think.

OpenStudy (dayakar):

A = pi * d^2/4 = pi* 40* 40 /4 = 400 8pi

OpenStudy (dayakar):

sorry A= pi* 400